Physics 152—Accelerated Physics II: Electricity, Magnetism, and Optics Homework Assignment #2 Revised Due Date: Wednesday, September 11, 2024, 9:30xam

Mastering Physics

Do the on-line assignment “HW #02.”

Pencil-and-Paper Problems

Ch.x22:

22.24, 22.27, and 22.28 (30 pts.xeach). For problem 22.28, note that you can’t directly use Gauss’s Law for points B and C because they are not symmetrically placed in the system. Instead you should use (but not re-derive) the result of Example 22.7 for a single surface and then superpose the results from all 4 surfaces together. You have to think carefully about the + or − for the field due to each surface. Drawing arrows on a figure is usually helpful.

Please write neatly and show your work clearly. I need to be able to follow your reasoning. Homework that is incomplete or difficult to understand will not get full credit.

Academic Honesty If you get bogged down with any of the problems, do not hesitate to discuss them with your instructor or with a fellow student. However, if you discuss a problem with anyone (besides your instructor) you should acknowledge that collaboration. Please see the Academic Honesty policy for more information about appropriate and inappropriate collaboration.